Perform three iteration of the bisection method compute the positive root of

the equation f(x) = x² – 5x + 2


1
Expert's answer
2021-06-09T17:46:43-0400
"f(x) = x^2 \u2013 5x + 2"

Initial values "a_0=0, b_0=1"


"f(a_0)=f(0)=(0)^2-5(0)+2=2"

"f(b_0)=f(1)=(1)^2-5(1)+2=-2"

"f(a_0)f(b_0)=2(-2)=-4<0"

"x_n=\dfrac{a_n+b_n}{2}"

"f(x_n)\leq\varepsilon=>answer=x_n"

"a_{n+1}=x_n, b_{n+1}=b_n, f(a_n)f(x_n)\geq 0"

"a_{n+1}=a_n, b_{n+1}=x_n, f(b_n)f(x_n)\geq 0"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 n & x & f(x) \\ \hline
 1 & 0.5 & -0.25 \\
 2 & 0.25 & 0.8125 \\
 3 & 0.375 & 0.265625 \\
 \hdashline
 
\end{array}"


Initial values "a_0=4, b_0=5"


"f(a_0)=f(4)=(4)^2-5(4)+2=-2"

"f(b_0)=f(5)=(5)^2-5(5)+2=2"

"f(a_0)f(b_0)=-2(2)=-4<0"

"x_n=\dfrac{a_n+b_n}{2}"

"f(x_n)\leq\varepsilon=>answer=x_n"

"a_{n+1}=x_n, b_{n+1}=b_n, f(a_n)f(x_n)\geq 0"

"a_{n+1}=a_n, b_{n+1}=x_n, f(b_n)f(x_n)\geq 0"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 n & x & f(x) \\ \hline
 1 & 4.5 & -0.25 \\
 2 & 4.75 & 0.8125 \\
 3 & 4.625 & 0.265625 \\
 \hdashline
 
\end{array}"